I am having a hard time figuring out this story problem: There are n teams and the number of games played is given by the polynomial n^2-n=N. A woman's softball league has 13 teams. What is the total number of games to be played?
n^2-13=N
I am not sure if I am writing out the equation correctly and where to go from there. Thank you for your time.

Answer by checkley71(8403)

WHY DO YOU REPLACE ONLY ONE OF THE n TERMS WITH 13?
YOU NEED TO REPLACE ALL n TERMS WITH 13 THUS:
n^2-n=N
13^2-13=N
N=169-13
N=156 IS THE NUMBER OF GAMES TO BE PLAYED BY THE 13 TEAMS.